please help with this factor problem.
Factor completely.
25x^2 - 9
thanks

Answer by prabhjyot(165)

25x^2 - 9=0
25 x^2 =9
x^2 =9/25
x = sqrt(9/25)
x=sqrt(0.36)
x=0.6